High Desert Potteryworks makes a variety of pottery products that it sells to retailers. The company uses a job-order costing system in which departmental predetermined overhead rates are used to apply manufacturing overhead cost to jobs. The predetermined overhead rate in the Molding Department is based on machine-hours, and the rate in the Painting Department is based on direct labor-hours. At the beginning of the year, the company provided the following estimates:

Department
Molding Painting
Direct labor-hours 34,000 53,900
Machine-hours 82,000 36,000
Fixed manufacturing overhead cost 180,400 485,100
Variable manufacturing overhead per machine-hour 3.00 -
Variable manufacturing overhead per direct labor-hour - 5.00

Job 205 was started on August 1 and completed on August 10. The company's cost records show the following information concerning the job:

Department
Molding Painting
Direct labor-hours 76 135
Machine-hours 360 73
Direct materials $ 934 $ 1,200
Direct labor cost $ 720 $ 960

Required:

1. Compute the predetermined overhead rates used in the Molding Department and the Painting Department.

2. Compute the total overhead cost applied to Job 205.

3-a. What would be the total manufacturing cost recorded for Job 205?

3-b. If the job contained 21 units, what would be the unit product cost?
Answer #1

1) Molding overhead rate = (180400/82000+3) = 5.20 per MH

Painting overhead rate = (485100/53900+5) = 14 per DLH

2) Total overhead cost applied = (360*5.2+135*14) = 3762

3) Total manufacturing cost

Direct material (934+1200) 2134
Direct labor (720+960) 1680
Overhead 3762
Total manufacturing cost 7576

4) Unit product cost = 7576/21 = 360.76